40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4987     Entry time: Tue Jul 19 09:19:14 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

The PRM watchdogs were tripped. The side was kicked up to 180mV Damping was restored.

ELOG V3.1.3-